# Local order metrics for many-particle systems across length scales

### v1

@@ Abstract
Formulating order metrics that sensitively quantify the degree of order/disorder in many-particle systems in $d$-dimensional Euclidean space $\mathbb{R}^{d}$ across length scales is an outstanding challenge in physics, chemistry, and materials science. Since an infinite set of $n$-particle correlation functions is required to fully characterize a system, one must settle for a reduced set of structural information, in practice. We initiate a program to use the number variance $\sigma_{N}^{2}(R)$ associated with a spherical sampling window of radius $R$ (which encodes pair correlations) and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$ that depends on two specified radial distances $R_{i}$ and $R_{j}$. Across the first three space dimensions ($d=1,2,3$), we find these metrics can sensitively describe and categorize the degree of order/disorder of 41 different models of antihyperuniform, nonhyperuniform, disordered hyperuniform, and ordered hyperuniform many-particle systems at a specified length scale $R$. Using our local variance metrics, we demonstrate the importance of assessing order/disorder with respect to a specific value of $R$. These order metrics could also aid in the inverse design of structures with prescribed scale-specific degrees of order/disorder that yield desired physical properties. In future work, it would be fruitful to explore the use of higher-order moments [S. Torquato et al., Phys. Rev. X, 11, 021028 (2021)] to devise even more sensitive order metrics.
@@ I Introduction
The classification of both ordered and disordered many-particle systems is a difficult and outstanding problem in physics, chemistry, and materials science. To fully characterize the microstructure of a many-particle system in $d$-dimensional Euclidean space $\mathbb{R}^{d}$, one requires an infinite set of $n$-particle probability density functions $\rho_{n}(\mathbf{r}^{n})$ associated with finding any $n$ particles in a configuration $\mathbf{r}^{n}$ [1, 2]. Such complete information is virtually never available in practice, so we must settle for reduced structural information, e.g., by only considering lower-order correlation functions. Subject to this necessary restriction, several scalar order metrics (see, e.g., [3, 4, 5, 6, 7, 8]) have been proposed that have been fruitfully applied to arrangements of nonoverlapping spheres (sphere packings) [4, 3, 9, 5, 10], simple liquids [11, 4, 12, 13], glasses [4, 14], water [15, 16, 17], disordered ground states [8, 18], random media [19, 20, 21, 22, 23], the prime numbers [24], two-phase media [25], and random fields [26].
@@
is the volume of a $d$-dimensional sphere of radius $R$ and $\Gamma(x)$ is the gamma function. In other words, the system is characterized by large-scale density fluctuations that are anomalously suppressed compare to those of a typical disordered system. Equivalently, a system is hyperuniform if $S(\mathbf{k})\rightarrow 0$ as $|\mathbf{k}|\rightarrow 0$.
@@
In this work, we initiate a program to use the scaled number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$ that depends on two specified radial distances $R_{i}$ and $R_{j}$ as an order metric for antihyperuniform, nonhyperuniform, disordered hyperuniform, and ordered hyperuniform point configurations at any length scale $R$. Specifically, at a specified $R$, lower (higher) number density fluctuations as measured by $\sigma_{N}^{2}(R)$ are used as a measure of a greater degree of order (disorder). Following Refs. 25 and 26, which profitably use a $1/v_{1}(R)$ scaling in the formulation of their $\sigma_{V}^{2}(R)$- and $\sigma_{F}^{2}(R)$-based order metrics, we demonstrate that $1/v_{1}(R)$ is also a good choice of scale for a $\sigma_{N}^{2}(R)$-based metric. Across the first three space dimensions ($d=1,2,$ and 3) we find that $\sigma_{N}^{2}(R)/v_{1}(R)$ and $\Sigma_{N}(R_{i},R_{j})$ can sensitively describe the degree of order/disorder of 41 different models of the aforementioned broad spectrum of many-particle systems at short, intermediate, and large length scales consistently with physical intuition. While many existing translational order metrics (see, e.g., Refs. 6, 7 and references therein) attempt to quantify the degree of order/disorder of a system globally (i.e., not at some selected length scale), our local variance metrics depend explicitly on a length scale $R$, and thus can sensitively and robustly characterize order/disorder on any prescribed length scale. We compare and contrast our number variance-based metrics and existing ones in Sec. V. Moreover, we demonstrate the length scale dependence of order/disorder rankings across models in a given space dimension, providing additional evidence that it is important to assess order/disorder with respect to a specific length scale [25, 26]. To aid in materials design, these order metrics could be applied to inverse techniques [47, 48, 49, 50] to generate structures with prescribed scale-specific degrees of order/disorder, yielding desired physical properties.
The rest of the paper is structured as follows. Section II contains pertinent background information and mathematical definitions. Section III describes the large variety of systems we apply our order metrics to. In Sec. IV we present our results, and in Section V we offer our conclusions and potential areas of future study.
@@ II.1 Pair statistics of many-particle systems
The structure factor $S(k)$ is related to the Fourier transform of $h(\mathbf{r})$ ($\tilde{h}(\mathbf{r})$) via
@@
Figure 1: Schematics indicating a circular observation window $\Omega$ of radius $R$ in two dimensions and its centroid $x_{0}$ for disordered nonhyperuniform (left), ordered hyperuniform (center), and disordered hyperuniform (right) point patterns. In each example, the number of points within the window $\Omega$ will fluctuate as the window position $x_{0}$ varies. The number variance $\sigma_{N}^{2}(R)$ will asymptotically scale like $R^{2}$ for the nonhyperuniform point pattern (right) and like $R$ for the ordered (center) and disordered (right) hyperuniform patterns; see Sec. IIB. for the other possible large-$R$ scalings.
@@ II.2 Number variance and hyperuniformity of many-particle systems
is the intersection volume of two $d$-dimensional spherical windows of radius R, scaled by the volume of a $d$-dimensional sphere $v_{1}(R)$ [given by Eq. (2)] whose centers are separated by a distance $r$ [9, 51], and $\tilde{\alpha}_{2}(k;R)$ is its Fourier transform. Figure 1 depicts schematically how one can compute the number variance $\sigma_{N}^{2}(R)$ directly from general point patterns by window sampling at some fixed radius $R$. We describe the computation of $\sigma_{N}^{2}(R)$ via direct sampling of a point pattern in more detail in Appendix A.
@@
The hyperuniformity exponent $\alpha$ (see Eq. (3)), can be used to divide hyperuniform systems into three distinct classes that describe their associated large-$R$ behaviors of $\sigma_{N}^{2}(R)$ [9, 42]:
@@
For a “typical” nonhyperuniform system, $S(0)$ is bounded [42], and for antihyperuniform systems, $S(0)$ is unbounded, i.e.,
@@ II.3 Translational order metrics
Naturally, there is an enormous family of scalar functions that depend on the particle positions that could in principle be used as an order metric, but they are not necessarily good order metrics. Based on experiences with existing order metrics, Kansal et al. [5] suggested that a good order metric is one that:
  1. 1.
    is sensitive to any type of ordering in a system and should not be biased toward any reference system;
  2. 2.
    reflects the hierarchy of ordering between prototypical systems given by physical intuition;
  3. 3.
    detects order at any length scale;
  4. 4.
    is affected by the variety of local configurational patterns and the spatial distribution of these patterns;
@@
The order metric $T$ measures the degree of translational order in a system of interest relative to the perfect face-centered cubic (FCC) lattice structure at the same number density [36, 4]. Specifically,
@@
where $n_{i}$ (for the system of interest) indicates the average occupation number for the thin spherical shell located a distance away from a reference particle equal to the $i$th nearest-neighbor separation for the FCC lattice at that number density, and $N_{C}$ is the total number of shells considered. Similarly, $n_{i}^{ideal}$ and $n_{i}^{FCC}$ are the corresponding shell occupation numbers for the Poisson point process and the FCC lattice, respectively. $T$ notably does not meet the aforementioned property 1 of a good order metric because it explicitly requires comparison to the FCC system.
@@
where $\xi=r\rho^{1/d}$, and $\xi_{C}$ is some cutoff distance. Density-density correlations are detected by integrating over the absolute value of the total correlation function. We note also that $T^{*}$ was designed to assess the degree of order/disorder globally for sphere packings specifically, and thus, its application is limited in scope.
@@
where $k_{B}$ is Boltzmann’s constant. This metric is essentially the multiparticle correlation function expansion of the excess entropy—relative to an ideal gas at the same number density—truncated at pair statistics. Specifically, the positive semi-definite quantity $-s^{(2)}/k$ is used as an order metric.
@@
Similarly to $T^{*}$, density-density correlations are detected by integrating over the total correlation function, but with a different weighting. Such a metric is advantageous over those mentioned previously because it can be formulated in terms of direct space or reciprocal space pair statistics. In addition, a local version of $\tau$ in reciprocal space, denoted by $\tau(K)$ has been formulated by introducing an upper bound on the integration [20]:
@@ III Hyperuniform and nonhyperuniform Models
We consider 12 different models of statistically homogeneous many-particle systems in one, two, and three dimensions: eight nonhyperuniform models, two of which are antihyperuniform, and four hyperuniform models. Our nonhyperuniform models are chosen such that they span a representative subset of the possible range of $S(0)$ values and our hyperuniform models are chosen such that they span disordered, ordered, stealthy (defined below), and nonstealthy varieties. A set of figures depicting representative configurations and pair statistics for each model for $d=1,2,3$ is given in the Supplemental Material [56]. To compare each model in a given space dimension consistently, we scale all distances by $\rho^{1/d}$, which, unless stated otherwise, is a reasonable estimate of the mean nearest-neighbor distance in the system.
@@ III.1.1 Periodic systems
where $\mathcal{L}^{*}$ denotes the reciprocal lattice of $\mathcal{L}$ and $\delta(x)$ is the Dirac delta function. The structure factor of all Bravais lattices for all wavenumbers up to the first Bragg peak is identically 0, which is referred to as stealthy hyperuniformity and makes them class I hyperuniform with $\sigma_{N}^{2}(R)\sim R^{d-1}$ scaling in the large-$R$ limit [57, 58, 8, 42]. Specifically, we examine the 1,2 and 3-dimensional (1D, 2D, and 3D) hypercubic lattices $\mathbb{Z}^{d}$, the 2D and 3D root lattices $A_{d}$ ($A_{3}$ is also known as the face-centered-cubic (FCC) lattice) and the dual lattice to the FCC lattice $A_{3}^{*}$ (or the body-centered-cubic (BCC) lattice). Here, we use Eqs. (23) and (9) to compute the number variance for the set of lattices listed above.
@@ III.1.2 Quasiperiodic systems
We also consider 1D and 2D quasiperiodic many-particle systems that have long-ranged orientational order, but exhibit quasiperiodic rather than periodic translational order [59]. The structure factors of such systems is composed of a dense set of Bragg peaks separated by gaps of arbitrarily small size [60]. For $d=1$, we examine the Fibonacci quasiperiodic chain, which has $\alpha=3$ [61] and for $d=2$, we examine the quasiperiodic Penrose tiling, which has $\alpha=6$ [62, 63]. Both these models are class I hyperuniform and thus have $\sigma_{N}^{2}(R)\sim R^{d-1}$ scaling in the large-$R$ limit. Here, for the 1D Fibonacci and 2D Penrose quasicperiodic systems, we directly sample the number variance using configurations generated via the substitution tiling method with $N=832040$ particles described in Ref. [64] and generated from the generalized dual method with $N=167761$ described in Refs. 65, 66, 67, respectively.
@@ III.1.3 Disordered stealthy hyperuniform systems
Disordered stealthy hyperuniform many-particle systems are class I hyperuniform with $\sigma_{N}^{2}(R)\sim R^{d-1}$ scaling in the large-$R$ limit and have structure factors that vanish in a spherical region around the origin, i.e., $S(\mathbf{k})=0$ for $0<\mathbf{k}\leq K$, but, unlike periodic systems, are isotropic and do not have Bragg peaks. Such systems can be generated using the collective-coordinate optimization scheme that involves finding the highly degenerate ground states of a class of bounded pair potentials with compact support in Fourier space, which are stealthy and hyperuniform by construction [68, 57, 69, 58, 8, 70, 42, 18]. The parameter $\chi$ is a dimensionless measure of the ratio of constrained degrees of freedom (i.e., the number of wave vectors contained within the cutoff wavenumber $K$) to the total degrees of freedom (roughly $dN$, where $N$ is the total number of points in the system). Many-particle systems with small $\chi$ (relatively unconstrained) will disordered, and as $\chi$ increases, the degree of short-range order increases within the “disordered regime” ($\chi<1/2$ for $d=2$ and 3 [8]). For $d=1$, the disordered regime only extends to $\chi<1/3$ [70]. Here, we directly sample the number variance for “entropically favored” disordered stealthy hyperuniform many-particle systems generated via the procedure in Ref. 70. In particular, we examine “high-$\chi$” stealthy hyperuniform systems, i.e., $\chi=0.3$ for $d=1$ and $\chi=0.49$ for $d=2$ and 3. For $d=1$, we consider 900 configurations with $N=1000$, for $d=2$, we consider 700 configurations with $N=10^{4}$, and for $d=3$, we consider 1000 configurations with $N=8000$.

### v2

@@ Abstract
Formulating order metrics that sensitively quantify the degree of order/disorder in many-particle systems in $d$-dimensional Euclidean space $\mathbb{R}^{d}$ across length scales is an outstanding challenge in physics, chemistry, and materials science. Since an infinite set of $n$-particle correlation functions is required to fully characterize a system, one must settle for a reduced set of structural information, in practice. We initiate a program to use the local number variance $\sigma_{N}^{2}(R)$ associated with a spherical sampling window of radius $R$ (which encodes pair correlations) and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$ that depends on two specified radial distances $R_{i}$ and $R_{j}$. Across the first three space dimensions ($d=1,2,3$), we find these metrics can sensitively describe and categorize the degree of order/disorder of 41 different models of antihyperuniform, nonhyperuniform, disordered hyperuniform, and ordered hyperuniform many-particle systems at a specified length scale $R$. Using our local variance metrics, we demonstrate the importance of assessing order/disorder with respect to a specific value of $R$. These local order metrics could also aid in the inverse design of structures with prescribed length-scale-specific degrees of order/disorder that yield desired physical properties. In future work, it would be fruitful to explore the use of higher-order moments of the number of points within a spherical window of radius $R$ [S. Torquato et al., Phys. Rev. X, 11, 021028 (2021)] to devise even more sensitive order metrics.
@@ I Introduction
The classification of both ordered and disordered many-particle systems is a difficult and outstanding problem in physics, chemistry, and materials science. To fully characterize the microstructure of a many-particle system in $d$-dimensional Euclidean space $\mathbb{R}^{d}$, one requires an infinite set of $n$-particle probability density functions $\rho_{n}(\mathbf{r}^{n})$ associated with finding any $n$ particles in a configuration $\mathbf{r}^{n}$ [1, 2]. Such complete information is virtually never available in practice, so we must settle for reduced structural information, e.g., by only considering lower-order correlation functions. Subject to this necessary restriction, several scalar order metrics (see, e.g., Refs. [3, 4, 5, 6, 7, 8]) have been proposed that have been fruitfully applied to arrangements of nonoverlapping spheres (sphere packings) [4, 3, 9, 5, 10], simple liquids [11, 4, 12, 13], glasses [4, 14], water [15, 16, 17], disordered ground states [8, 18], random media [19, 20, 21, 22, 23], the prime numbers [24], two-phase media [25], and random fields [26].
@@
is the volume of a $d$-dimensional sphere of radius $R$ and $\Gamma(x)$ is the gamma function. In other words, the system is characterized by large-scale density fluctuations that are anomalously suppressed compared to those of a typical disordered system. Equivalently, a system is hyperuniform if $S(\mathbf{k})\rightarrow 0$ as $|\mathbf{k}|\rightarrow 0$.
@@
In this work, we initiate a program to use the scaled local number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$ that depends on two specified radial distances $R_{i}$ and $R_{j}$ as an order metric for antihyperuniform, nonhyperuniform, disordered hyperuniform, and ordered hyperuniform point configurations at any length scale $R$. Specifically, at a specified $R$, lower (higher) number density fluctuations as measured by $\sigma_{N}^{2}(R)$ are used as a measure of a greater degree of order (disorder). Following Refs. 25 and 26, which profitably use a $1/v_{1}(R)$ scaling in the formulation of their $\sigma_{V}^{2}(R)$- and $\sigma_{F}^{2}(R)$-based order metrics, we demonstrate that $1/v_{1}(R)$ is also a good choice of scale for a $\sigma_{N}^{2}(R)$-based metric. Across the first three space dimensions ($d=1,2,$ and 3) we find that $\sigma_{N}^{2}(R)/v_{1}(R)$ and $\Sigma_{N}(R_{i},R_{j})$ can sensitively describe the degree of order/disorder of 41 different models of the aforementioned broad spectrum of many-particle systems at short, intermediate, and large length scales consistently with physical intuition. While many existing translational order metrics (see, e.g., Refs. 6, 7 and references therein) attempt to quantify the degree of order/disorder of a system globally (i.e., not at some selected length scale), our local variance metrics depend explicitly on a length scale $R$, and thus can sensitively and robustly characterize order/disorder on any prescribed length scale. We compare and contrast our number variance-based metrics and existing ones in Sec. V. Moreover, we demonstrate the length scale dependence of order/disorder rankings across models in a given space dimension, providing additional evidence that it is important to assess order/disorder with respect to a specific length scale [25, 26]. To aid in materials design, these order metrics could be applied to inverse techniques [47, 48, 49, 50] to generate structures with prescribed scale-specific degrees of order/disorder, yielding desired physical properties.
The rest of the paper is structured as follows. Section II contains pertinent background information and mathematical definitions. Section III describes the large variety of systems to which we apply our order metrics. In Sec. IV we present our results, and in Sec. V we offer our conclusions and potential areas for future study.
@@ II.1 Pair statistics of many-particle systems
The structure factor $S(k)$ is related to the Fourier transform of $h(\mathbf{r})$ [$\tilde{h}(\mathbf{r})$] via
@@
Figure 1: Schematics indicating a circular observation window $\Omega$ of radius $R$ in two dimensions and its centroid $x_{0}$ for disordered nonhyperuniform (left), ordered hyperuniform (center), and disordered hyperuniform (right) point patterns. In each example, the number of points within the window $\Omega$ will fluctuate as the window position $x_{0}$ varies. The number variance $\sigma_{N}^{2}(R)$ will asymptotically scale like $R^{2}$ for the nonhyperuniform point pattern (right) and like $R$ for the ordered (center) and disordered (right) hyperuniform patterns; see Sec. II B for the other possible large-$R$ scalings.
@@ II.2 Number variance and hyperuniformity of many-particle systems
is the intersection volume of two $d$-dimensional spherical windows of radius R, scaled by the volume of a $d$-dimensional sphere $v_{1}(R)$ [given by Eq. (2)] whose centers are separated by a distance $r$ [9, 51], and $\tilde{\alpha}_{2}(k;R)$ is its Fourier transform. Figure 1 depicts schematically how one can compute the number variance $\sigma_{N}^{2}(R)$ directly from general point patterns by window sampling at some fixed radius $R$. We describe the computation of $\sigma_{N}^{2}(R)$ via direct sampling of a point pattern in more detail in the Appendix.
@@
The hyperuniformity exponent $\alpha$ [see Eq. (3)], can be used to divide hyperuniform systems into three distinct classes that describe their associated large-$R$ behaviors of $\sigma_{N}^{2}(R)$ [9, 42]:
@@
For a typical nonhyperuniform system, $S(0)$ is bounded [42], and for antihyperuniform systems, $S(0)$ is unbounded, i.e.,
@@
In summary, the scalings (15) and (16) provide a complete characterization of large-scale density fluctuations that span all configurations, hyperuniform or not or ordered or not [46].
@@ II.3 Translational order metrics
Naturally, there is an enormous family of scalar functions that depend on the particle positions that could in principle be used as an order metric, but they are not necessarily good order metrics. Based on experiences with existing order metrics, Kansal et al. [5] suggested that a good order metric is one that (1) is sensitive to any type of ordering in a system and should not be biased toward any reference system, (2) reflects the hierarchy of ordering between prototypical systems given by physical intuition, (3) detects order at any length scale, and (4) is affected by the variety of local configurational patterns and the spatial distribution of these patterns.
@@
The order metric $T$ measures the degree of translational order in a system of interest relative to the perfect face-centered-cubic (fcc) lattice structure at the same number density [36, 4]. Specifically,
@@
where $n_{i}$ (for the system of interest) indicates the average occupation number for the thin spherical shell located a distance away from a reference particle equal to the $i$th nearest-neighbor separation for the fcc lattice at that number density and $N_{C}$ is the total number of shells considered. Similarly, $n_{i}^{ideal}$ and $n_{i}^{fcc}$ are the corresponding shell occupation numbers for the Poisson point process and the fcc lattice, respectively. $T$ notably does not meet the aforementioned property (1) of a good order metric because it explicitly requires comparison to the fcc system.
@@
where $\xi=r\rho^{1/d}$ and $\xi_{C}$ is some cutoff distance. Density-density correlations are detected by integrating over the absolute value of the total correlation function. We note also that $T^{*}$ was designed to assess the degree of order/disorder globally for sphere packings specifically, and thus, its application is limited in scope.
@@
where $k_{B}$ is Boltzmann’s constant. This metric is essentially the multiparticle correlation function expansion of the excess entropy—relative to an ideal gas at the same number density—truncated at pair statistics. Specifically, the positive semidefinite quantity $-s^{(2)}/k$ is used as an order metric.
@@
Similarly to $T^{*}$, density-density correlations are detected by integrating over the total correlation function, but with a different weighting. Such a metric is advantageous over those mentioned previously because it can be formulated in terms of direct space or reciprocal space pair statistics. In addition, a local version of $\tau$ in reciprocal space, denoted by $\tau(K)$, has been formulated by introducing an upper bound on the integration [20]:
@@ III Hyperuniform and nonhyperuniform Models
We consider 12 different models of statistically homogeneous (translationally invariant) many-particle systems in one, two, and three dimensions: eight nonhyperuniform models, two of which are antihyperuniform, and four hyperuniform models. Our nonhyperuniform models are chosen such that they span a representative subset of the possible range of $S(0)$ values and our hyperuniform models are chosen such that they span disordered, ordered, stealthy (defined below), and nonstealthy varieties. A set of figures depicting representative configurations and pair statistics for each model for $d=1,2,3$ is given in the Supplemental Material [56]. To compare each model in a given space dimension consistently, we scale all distances by $\rho^{1/d}$, which, unless stated otherwise, is a reasonable estimate of the mean nearest-neighbor distance in the system.
@@ III.1.1 Periodic systems
where $\mathcal{L}^{*}$ denotes the reciprocal lattice of $\mathcal{L}$ and $\delta(x)$ is the Dirac delta function. The structure factor of all Bravais lattices for all wavenumbers up to the first Bragg peak is identically 0, which is referred to as stealthy hyperuniformity and makes them class I hyperuniform with $\sigma_{N}^{2}(R)\sim R^{d-1}$ scaling in the large-$R$ limit [57, 58, 8, 42]. Specifically, we examine the one, two and three-dimensional (1D, 2D, and 3D) hypercubic lattices $\mathbb{Z}^{d}$, the 2D and 3D root lattices $A_{d}$ [$A_{3}$ is also known as the face-centered-cubic (fcc) lattice] and the dual lattice to the fcc lattice $A_{3}^{*}$ [or the body-centered-cubic (bcc) lattice]. Here, we use Eqs. (23) and (9) to compute the number variance for the set of lattices listed above. While periodic systems are not statistically homogeneous, the exact calculations of their number variances only involve their radial statistics [1, 42].
@@ III.1.2 Quasiperiodic systems
We also consider 1D and 2D quasiperiodic many-particle systems that have long-ranged orientational order, but exhibit quasiperiodic rather than periodic translational order [59]. The structure factors of such systems are composed of a dense set of Bragg peaks separated by gaps of arbitrarily small size [60]. For $d=1$, we examine the Fibonacci quasiperiodic chain, which has $\alpha=3$ [61] and for $d=2$, we examine the quasiperiodic Penrose tiling, which has $\alpha=6$ [62, 63]. Both these models are class I hyperuniform and thus have $\sigma_{N}^{2}(R)\sim R^{d-1}$ scaling in the large-$R$ limit. Here, for the 1D Fibonacci and 2D Penrose quasicperiodic systems, we directly sample the number variance using configurations generated via the substitution tiling method with $N=832040$ particles described in Ref. [64] and generated from the generalized dual method with $N=167761$ described in Refs. 65, 66, 67, respectively.
@@ III.1.3 Disordered stealthy hyperuniform systems
Disordered stealthy hyperuniform many-particle systems are class I hyperuniform with $\sigma_{N}^{2}(R)\sim R^{d-1}$ scaling in the large-$R$ limit and have structure factors that vanish in a spherical region around the origin, i.e., $S(\mathbf{k})=0$ for $0<\mathbf{k}\leq K$, but, unlike periodic systems, are isotropic and do not have Bragg peaks. Such systems can be generated using the collective-coordinate optimization scheme that involves finding the highly degenerate ground states of a class of bounded pair potentials with compact support in Fourier space, which are stealthy and hyperuniform by construction [68, 57, 69, 58, 8, 70, 42, 18]. The parameter $\chi$ is a dimensionless measure of the ratio of constrained degrees of freedom (i.e., the number of wave vectors contained within the cutoff wavenumber $K$) to the total degrees of freedom (roughly $dN$, where $N$ is the total number of points in the system). Many-particle systems with small $\chi$ (relatively unconstrained) will have short-range disorder, and as $\chi$ increases, the degree of short-range order increases within the “disordered regime” ($\chi<1/2$ for $d=2$ and 3 [8]). For $d=1$, the disordered regime only extends to $\chi<1/3$ [70]. Here, we directly sample the number variance for “entropically favored” disordered stealthy hyperuniform many-particle systems generated via the procedure in Ref. 70. In particular, we examine “high-$\chi$” stealthy hyperuniform systems, i.e., $\chi=0.3$ for $d=1$ and $\chi=0.49$ for $d=2$ and 3. For $d=1$, we consider 900 configurations with $N=1000$, for $d=2$, we consider 700 configurations with $N=10^{4}$, and for $d=3$, we consider 1000 configurations with $N=8000$.